# Rigidity and Structural Asymmetry of Bounded Solutions

### 16 Mar 2026 v1

@@ 3 Proof of the Theorem 4
$I(t)=\int_{1}^{t}\left[\frac{\left(\frac{t}{v}\right)^{\sigma}-\left(\frac{t}{v}\right)^{1-\sigma}}{\Im\big(\delta_{\eta,1+i\Im(s)}(t)\big)}\right]v^{-1}\,\Im\big(\delta_{\eta,1+i\Im(s)}(v)\big)\,dv$
From estimate (9), as $t\to+\infty$ we have
$\Im\big(\delta_{\eta,1+i\Im(s)}(t)\big)=-\,t\Im\big(\mu_{\eta}(1+i\tau)\big)+O(1).$
Hence
$\frac{\Im\big(\delta_{\eta,1+i\tau}(v)\big)}{\Im\big(\delta_{\eta,1+i\tau}(t)\big)}=\frac{v}{t}+o(1)$
uniformly for $1\leq v\leq t$. Substituting this relation into the integral kernel gives
$I(t)\sim\int_{1}^{t}\left[t^{\sigma}v^{-\sigma}-t^{1-\sigma}v^{\sigma-1}\right]v^{-1}\frac{v}{t}\,dv.$
Thus
$I(t)\sim\frac{1}{t}\int_{1}^{t}\left(t^{\sigma}v^{-\sigma}-t^{1-\sigma}v^{\sigma-1}\right)dv.$
We compute
$\int_{1}^{t}t^{\sigma}v^{-\sigma}dv=t^{\sigma}\frac{t^{1-\sigma}-1}{1-\sigma}=\frac{t-t^{\sigma}}{1-\sigma},$
@@
$\int_{1}^{t}t^{1-\sigma}v^{\sigma-1}dv=t^{1-\sigma}\frac{t^{\sigma}-1}{\sigma}=\frac{t-t^{1-\sigma}}{\sigma}.$
Therefore
$I(t)\sim\frac{1}{t}\left[\frac{t-t^{\sigma}}{1-\sigma}-\frac{t-t^{1-\sigma}}{\sigma}\right].$
After simplification we obtain
$I(t)\sim\frac{(2\sigma-1)t-\sigma t^{\sigma}+(1-\sigma)t^{1-\sigma}}{\sigma(1-\sigma)t}.$
Substituting into the previous expression yields
$\frac{\Im\big(\delta_{\eta,s}(t)\big)}{\Im\big(\delta_{\eta,1+i\tau}(t)\big)}\sim 1-\frac{\sigma(1-\sigma)}{2\sigma-1}\frac{(2\sigma-1)t-\sigma t^{\sigma}+(1-\sigma)t^{1-\sigma}}{\sigma(1-\sigma)t}.$
After simplification we obtain
$\frac{\Im\big(\delta_{\eta,s}(t)\big)}{\Im\big(\delta_{\eta,1+i\tau}(t)\big)}\sim\frac{\sigma}{(2\sigma-1)t^{1-\sigma}}-\frac{1-\sigma}{(2\sigma-1)t^{\sigma}}>0\quad\text{as}\quad t\to+\infty.$
If $\mu_{\eta}(s)=\mu_{\eta}(1-\overline{s})=0$, Lemma 3 implies that there exists $C_{s}>0$ such that
$\left|\Im\big(\delta_{\eta,s}(t)\big)-\Im\left(\rho\left(\frac{s-1}{s}+\frac{\overline{s}}{1-\overline{s}}\right)t^{-i\Im(s)}\right)\right|<C_{s}t^{-1},\quad\forall t\geq 1.$
Since $s\in B$, the $\Re(s)\neq\frac{1}{2}$ and $\Im(s)\neq 0$. Use the fact that $\rho\neq 0$, then
$\rho\big(\frac{s-1}{s}+\frac{\overline{s}}{1-\overline{s}}\big)\neq 0.$
Hence, there exists a sequence $(t_{n})_{n\in\mathbb{N}}\subset[1,+\infty)$ such that $\lim_{n\to+\infty}t_{n}=+\infty$ and such that
$\Im\big(\delta_{\eta,s}(t_{n})\big)=0,\quad\forall n\in\mathbb{N},$
From equation (9) and the assumption $\Im(\mu_{\eta}(1+i\beta))\neq 0$, there exists $n_{s}\in\mathbb{N}$ such that
$\frac{\Im\big(\delta_{\eta,s}(t_{n})\big)}{\Im\big(\delta_{\eta,1+i\Im(s)}(t_{n})\big)}=0,\quad\forall n\geq n_{s},$
In other words, the function
$t\mapsto\frac{\Im\big(\delta_{\eta,s}(t)\big)}{\Im\big(\delta_{\eta,1+i\Im(s)}(t)\big)}$
oscillates like $t^{-1}\Im\big(\omega_{s}t^{-i\Im(s)}\big)$ with $\omega_{s}\in\mathbb{C}^{*}$, preventing the ratio in (10) from remaining strictly positive or converging as required, thus yielding a contradiction. ∎

### 15 Apr 2026 v3

@@ 4 Proof of the Theorem 4
$K(\sigma,c,t)$ $:=\int_{1}^{t}\mathcal{A}_{c}(u)u^{-\sigma}\int_{1}^{u}\,\frac{v^{-1+\sigma}}{\mathcal{A}_{c}(v)}\,dv\,du$
$-\int_{1}^{t}\mathcal{A}_{c}(u)u^{-\sigma}\int_{1}^{u}\,v^{-2(1-\sigma)}\int_{1}^{v}z\Big(\frac{d}{dz}\frac{z^{-\sigma}}{\mathcal{A}_{c}(z)}\Big)\,\,dz\,dv\,du,\quad\forall t\geq 1.$
We prove the theorem by contradiction. Suppose that $\mu_{\eta}(s)=\mu_{\eta}(1-\overline{s})=0$. Using the notation of $\delta_{\eta,s}$ given in Notation 6,it follows from the second item of the Lemma 3 that there exists a constant $C_{s}>0$ such that such that
$\left|\delta_{\eta,s}(t)\right|<C_{s},\quad\forall t\geq 1.$
Using the following compact convergence
$\lim_{c\to+\infty}c\,\mathcal{A}_{c}(u)=\frac{1}{\pi},$
We obtain
$\lim_{c\to+\infty}\Big|t\Phi_{c}[{\delta}_{\eta,s}](t)\Big|=\lim_{c\to+\infty}\Big|c\int_{1}^{t}\,\mathcal{A}_{c}(u)u^{-1}\delta_{\eta,s}(u)\,du\Big|$ $\leq\frac{1}{\pi}\,C_{s}\,\ln(t),\quad\forall t\geq 1.$
By the notation of $\tilde{\delta}_{s}$ in equation (10), we get
$\lim_{c\to+\infty}\Big|t\Phi_{c}[\tilde{\delta}_{s}](t)\Big|<\frac{1}{\pi}|\omega_{\tau}|\,C_{s}\,\ln(t),\quad\forall t\geq 1.$
Equation (15) gives
$\lim_{c\to+\infty}\frac{\Big|\Phi_{c}[\tilde{\delta}_{1+i\tau}](t)\Big|}{\ln(t)}\Big(t-\sigma(1-\sigma)K(\sigma,c,t)\Big)<\frac{1}{\pi}|\omega_{\tau}|\,C_{s},\quad\quad\forall t\geq t_{c_{s}}.$ (16)
Now, let us examine the asymptotic behavior of $K(\sigma,c,t)$. We recall the definition
$\mathcal{A}_{c}(t)=\frac{1}{2}+\frac{1}{\pi}\arctan(t-c),\qquad t\geq 1,$
@@
$K(\sigma,c,t)$ $=\int_{1}^{t}\mathcal{A}_{c}(u)u^{-\sigma}\int_{1}^{u}\,\frac{v^{-1+\sigma}}{\mathcal{A}_{c}(v)}\,dv\,du$
$-\int_{1}^{t}\mathcal{A}_{c}(u)u^{-\sigma}\int_{1}^{u}\,v^{-2(1-\sigma)}\int_{1}^{v}z\Big(\frac{d}{dz}\frac{z^{-\sigma}}{\mathcal{A}_{c}(z)}\Big)\,\,dz\,dv\,du,\quad\forall t\geq 1.$
Use the integration by parts formula to obtain
$K(\sigma,c,t)=\frac{1}{2\sigma-1}\int_{1}^{t}\mathcal{A}_{c}(u)u^{-\sigma}W(u)\,du,$ (17)
with
$W(u):=\dfrac{1}{\mathcal{A}_{c}(1)}(u^{-1+2\sigma}-1)+u^{-1+2\sigma}\int_{1}^{u}\dfrac{v^{-\sigma}}{\mathcal{A}_{c}(v)}\,dv-\int_{1}^{u}\frac{v^{-1+\sigma}}{\mathcal{A}_{c}(v)}\,dv.$
Use the integration by part formula, to get
$\mathcal{A}_{c}(u)u^{-\sigma}W(u)$ $=\frac{1}{1-\sigma}-\frac{1}{\sigma}$
$-\frac{\mathcal{A}_{c}(u)}{1-\sigma}u^{-1+\sigma}\Big[\dfrac{\sigma}{\mathcal{A}_{c}(1)}+\int_{1}^{u}v^{1-\sigma}\frac{d}{dv}\dfrac{1}{\mathcal{A}_{c}(v)}\,dv\Big]$
$+\frac{\mathcal{A}_{c}(u)}{\sigma}u^{-\sigma}\Big[\dfrac{1-\sigma}{\mathcal{A}_{c}(1)}+\int_{1}^{u}v^{\sigma}\frac{d}{dv}\dfrac{1}{\mathcal{A}_{c}(v)}\,dv\Big]$
Since
$\frac{d}{du}\mathcal{A}_{c}(u)=\frac{1}{\pi}\frac{1}{1+(u-c)^{2}},\quad\forall u\geq 1,$
It follows that
$\mathcal{A}_{c}(u)u^{-\sigma}W(u)$ $=\frac{1}{1-\sigma}-\frac{1}{\sigma}$
$-\frac{\mathcal{A}_{c}(u)}{1-\sigma}u^{-1+\sigma}\Big[\dfrac{\sigma}{\mathcal{A}_{c}(1)}-\frac{1}{\pi}\int_{1}^{u}\dfrac{1}{(\mathcal{A}_{c}(v))^{2}}\frac{v^{1-\sigma}}{1+(v-c)^{2}}\,dv\Big]$
$+\frac{\mathcal{A}_{c}(u)}{\sigma}u^{-\sigma}\Big[\dfrac{1-\sigma}{\mathcal{A}_{c}(1)}-\frac{1}{\pi}\int_{1}^{u}\dfrac{1}{(\mathcal{A}_{c}(v))^{2}}\frac{v^{\sigma}}{1+(v-c)^{2}}\,dv\Big]$
Then
$\lim_{c\to+\infty}\mathcal{A}_{c}(u)u^{-\sigma}W(u)=\frac{2\sigma-1}{\sigma(1-\sigma)}-\frac{\sigma}{1-\sigma}u^{-1+\sigma}+\frac{1-\sigma}{\sigma}u^{-\sigma}$
Inserting these into the integral for $K(\sigma,c,t)$ given by equation (17), we obtain the following equation
$\lim_{c\to+\infty}\Big(t-\sigma(1-\sigma)K(\sigma,c,t)\Big)=\frac{1}{2\sigma-1}\,\Big(\sigma t^{\sigma}-(1-\sigma)t^{1-\sigma}\Big).$
Using the following compact convergence
$\lim_{c\to+\infty}c\,\mathcal{A}_{c}(u)=\frac{1}{\pi},$
Thanks to equation (11), we get
$\lim_{t\to+\infty}\lim_{c\to+\infty}\Big|\Phi_{c}[\tilde{\delta}_{1+i\tau}](t)\Big|=\frac{1}{\pi}\big|\Re\big(\omega_{\tau}\alpha_{\tau})\big|>0.$
Since $\sigma\in(0,1)\setminus\frac{1}{2}$, we obtain
$\lim_{t\to+\infty}\lim_{c\to+\infty}\frac{\Big|\Phi_{c}[\tilde{\delta}_{1+i\tau}](t)\Big|}{\ln(t)}\Big(t-\sigma(1-\sigma)K(\sigma,c,t)\Big)=+\infty.$
We obtain a contradiction with equation (16). ∎